OpenStudy (anonymous):

4m^7/20n divided by 8m^4n/25 divide the write the quotient in lowest terms

OpenStudy (campbell_st):

well it looks like this \[\frac{4m^7}{20n} \div \frac{8m^4n}{25}\] the rule for dividing by a fraction is find the reciprocal of the 2nd fraction and multiply. or flip the 2nd and multiply so you get \[\frac{4m^7}{20n} \times \frac{25}{8m^4n}\] find the common factors that cancel before multiplying... then it shoule be easier to get an answer.